Given: fair coin, P(H)=P(T)=21
n=number of tosses,P(at least 2 heads)≥0.96
To find: minimum n with P(X≥2)≥0.96
P(X≥2)=1−P(X=0)−P(X=1)
Require P(X≥2)≥0.96
P(X=0)=(21)n
P(X=1)=n(21)n
P(X=0)+P(X=1)=2nn+1
P(X≥2)=1−2nn+1
1−2nn+1≥0.96⟺2nn+1≤0.04
Need 2nn+1≤0.04
n=7: 1288=161=0.0625>0.04 ✗
n=8: 2569=0.0352<0.04 ✓
P(X≥2)=1−2569=256247≈0.9648≥0.96 ✓
Minimum number of tosses=8